27. Determine the magnitude and direction of the force between two parallel wires 25m long and 4.0cm apart. each carrying 25 A in the same direction.
Answer #1

Length = L = 25 m

Current = I = 25 A

Separation = d = 4 cm

Force = F = (uo*I1*I2*L)/(2pi*d) = (2*10^-7*25*25*25)/(0.04) = 0.078125 N

As current flowing in same direction, then the force will be in opposite direction and hence will be attracting force.